# A Minimal Axio-dilaton Dark Sector

### v1

@@ 1.2 The Dark Axio-dilaton
Part of what makes this proposal attractive is the observation that the axion and dilaton have separately already been suggested as a Dark Matter or Dark Energy candidate. On one hand, axions have long been a leading candidate for Dark Matter [29, 30, 31] (for a review see [32]) because coherent field oscillations have an energy that falls within an expanding universe in the same was as does a nonrelativistic species of particle. The attractiveness of axions as dark-matter candidates has only grown as the viability of WIMP Dark Matter has decreased.
@@
The precise question we here ask is: can the axio-dilaton provide the one-stop shopping cosmologists need to describe both Dark Matter and Dark Energy? More specifically, recent studies show [35] that axio-dilaton models can produce acceptable cosmologies provided that Dark Matter is already present as a separate ingredient. We here ask whether the axion part of the axio-dilaton already used to describe Dark Energy can itself provide a minimal and predictive description of Dark Matter.
@@ 2.2.1 Adiabatic evolution
where $C$ is a proportionality constant (whose value is chosen to ensure the energy density of the axion fluid has the right size to be the Dark Matter). As expected, we find the standard $a^{-3}$ falloff for the energy of the fast oscillations, supplemented by the time-dependence of ${{\mathfrak{m}}}$.
@@ 3.1 CMB and power spectrum
Figure 1: Background and perturbative level plots of the axio-dilaton cosmology with a pure exponential dilaton potential. Top row shows the evolution of the relevant background energy densities when $\zeta=0.1$. Middle row shows the evolution of the dilaton field and associated baryon masses for a range of $\zeta$. Bottom row shows the corresponding angular and matter power spectra with the $\Lambda\rm CDM$ best-fit in solid blue. In all cases $\mathbf{g}=-10^{-3}$.
@@ 4 Summary and Conclusions
Because of the couplings to matter, we find that the largest effect on the CMB anisotropy power spectrum is at low multipoles. The variation of baryon and Dark Matter masses causes the gravitational potentials to evolve even in the matter dominated epoch according to the modified Einstein equations (2.30), (2.31) and (2.32), causing a large contribution to the integrated Sachs–Wolfe effect. We also calculated the quantity $f\sigma_{8}$, where $f$ is the growth rate of density perturbations in matter and $\sigma_{8}$ the normalization of the matter power spectrum at $R=8h^{-1}$pc, as a function of redshift. We find that for a pure exponential potential, this quantity behaves qualitatively similarly as in the $\Lambda$CDM model, but is usually larger. This behaviour is similar to what has been found in other interacting dark energy models [58, 59, 60]. If the potential has a local minimum, the dilaton oscillates around this minimum. As a consequence, both the background evolution and the growth rate evolution is modified, resulting in oscillatory behaviour of the quantity $f\sigma_{8}$, for which the amplitude is determined by $\zeta$, as can be seen in Fig. 5. This behaviour is significantly different from other interacting Dark Energy models.
@@ Acknowledgements
We thank Elsa Teixeira for the useful numerical resources. This work evolved out of discussions at the Astroparticle Symposium at the Institut Pascal. CB, CvdB and ACD thank the Institut Pascal for their hospitality during the programme. MM is also grateful for the hospitality of Perimeter Institute where part of this work was carried out. AS is supported by the W.D. Collins Scholarship. CvdB is supported by the Lancaster–Sheffield Consortium for Fundamental Physics under STFC grant: ST/X000621/1. ACD is partially supported by the Science and Technology Facilities Council (STFC) through the STFC consolidated grant ST/T000694/1. CB’s research was partially supported by funds from the Natural Sciences and Engineering Research Council (NSERC) of Canada. MM is supported in parts by the Mid-Career Research Program (2019R1A2C2085023) through the National Research Foundation of Korea Research Grants. MM is also supported Kavli IPMU which was established by the World Premier International Research Center Initiative (WPI), MEXT, Japan. This work was also supported by a grant from the Simons Foundation (1034867, Dittrich). Research at the Perimeter Institute is supported in part by the Government of Canada through NSERC and by the Province of Ontario through MRI.

### v2

@@ 1.2 The Dark Axio-dilaton
Part of what makes this proposal attractive is the observation that the axion and dilaton have separately already been suggested as a Dark Matter or Dark Energy candidate. On one hand, axions have long been a leading candidate for Dark Matter [29, 30, 31] (for a review see [32]) because coherent field oscillations have an energy that falls within an expanding universe in the same way as does a nonrelativistic species of particle. The attractiveness of axions as dark-matter candidates has only grown as the viability of WIMP Dark Matter has decreased [33].
@@
The precise question we here ask is: can the axio-dilaton provide the one-stop shopping cosmologists need to describe both Dark Matter and Dark Energy? More specifically, recent studies show [36] that axio-dilaton models can produce acceptable cosmologies provided that Dark Matter is already present as a separate ingredient. We here ask whether the axion part of the axio-dilaton already used to describe Dark Energy can itself provide a minimal and predictive description of Dark Matter. To answer this question we build on the previous work in [36], where the axion in all cases but one was taken to be massless, by giving the axion part an appropriate quadratic scalar potential, as known to be compatible with scalar field dark matter models [32].
@@ 2.2.1 Adiabatic evolution
where $C$ is a proportionality constant (whose value is chosen to ensure the energy density of the axion fluid has the right size to be the Dark Matter [^5] In practice we set the constant $C$ by ensuring that the fractional share of Dark Matter today $\Omega_{\textnormal{ax}0}h^{2}=0.120$ agrees with the Planck 2018 best-fit value taken from [45], where $h=0.674$, and return to the issue of physical production mechanisms responsible for setting this constant in §2.4 . ). As expected, we find the standard $a^{-3}$ falloff for the energy of the fast oscillations, supplemented by the time-dependence of ${{\mathfrak{m}}}$.
@@ 3.1 CMB and power spectrum
Figure 1: Background and perturbative level plots of the axio-dilaton cosmology with a pure exponential dilaton potential using CLASS [39]. Top row shows the evolution of the relevant background energy densities when $\zeta=0.1$. Middle row shows the evolution of the dilaton field and associated baryon masses for a range of $\zeta$. Bottom row shows the corresponding angular and matter power spectra with the $\Lambda\rm CDM$ best-fit in solid blue. In all cases $\mathbf{g}=-10^{-3}$.
@@ 3.2 Structure growth
It is interesting to compare the modifications in structure growth for the case when the dilaton is trapped within its quadratic minimum with the effects on the CMB angular power spectrum in the bottom row of figure 3. The case $\zeta=0.1$ is enough to give large deviations in structure growth while producing minimal deviations from $\Lambda$CDM for the low $\ell$ regime in the angular power spectrum. A full data analysis to examine whether these deviations could fit multiple combinations of cosmological datasets is therefore of interest for these models. Future experiments sensitive to the integrated Sachs-Wolfe effect such as CMB cluster cross-correlations are predicted to be able to place tighter constraints on these deviations, as discussed in [59, 60] .
@@ 4 Summary and Conclusions
Because of the couplings to matter, we find that the largest effect on the CMB anisotropy power spectrum is at low multipoles. The variation of baryon and Dark Matter masses causes the gravitational potentials to evolve even in the matter dominated epoch according to the modified Einstein equations (2.30), (2.31) and (2.32), causing a large contribution to the integrated Sachs–Wolfe effect. Similar studies on the integrated Sachs-Wolfe effect in the presence of modified gravity have been performed in for example [62]. These studies show it is complicated to draw conclusions from these types of deviations when comparing with multiple cosmological datasets. In order to infer if these deviations from $\Lambda$CDM can describe the cosmology better, a full data analysis would be required, which we leave for future work.
We also calculated the quantity $f\sigma_{8}$, where $f$ is the growth rate of density perturbations in matter and $\sigma_{8}$ the normalization of the matter power spectrum at $R=8h^{-1}$pc, as a function of redshift. We find that for a pure exponential potential, this quantity behaves qualitatively similarly as in the $\Lambda$CDM model, but is usually larger. This behaviour is similar to what has been found in other interacting dark energy models [63, 64, 65]. If the potential has a local minimum, the dilaton oscillates around this minimum. As a consequence, both the background evolution and the growth rate evolution is modified, resulting in oscillatory behaviour of the quantity $f\sigma_{8}$, for which the amplitude is determined by $\zeta$, as can be seen in Fig. 5. This behaviour is significantly different from other interacting Dark Energy models.
@@
Phantom Note Added
We briefly comment here on the possibility of apparent phantom behaviour in models like this, in view of the DESI results [66] that came out several months after this paper was posted. The main observation is that the dilaton-dependence we find here for the Dark Matter mass means that the actual equation of state parameter for the dilaton,
$\omega_{\chi}(\chi)=\frac{\chi^{\prime 2}-2a^{2}V(\chi)}{\chi^{\prime 2}+2a^{2}V(\chi)}\geq-1\,,$ (4.1)
can differ from the effective Dark Energy equation of state parameter $\omega_{\chi\,\text{eff}}$ one would infer if one were to mistakenly try to interpret the dark sector as a quintessence field coupled to ordinary Dark Matter (such as is done when using the $w_{0}w_{a}$ model), along the lines discussed in e.g. [67, 68, 69, 70]. For the cases studied here the effective dark energy equation of state would be given by
$\omega_{\chi\,\text{eff}}=\frac{\omega_{\chi}(\chi)}{1+\left[e^{\zeta(\chi-\chi_{0})}-1\right]\frac{\rho_{\text{ax}0}}{a^{3}\rho_{\chi}}},$ (4.2)
where $\rho_{\text{ax}\,0}=Cm_{\mathfrak{a}}/W(\chi_{0})$ and $\chi_{0}$ are respectively the dark matter energy density and value of the dilaton today.
Figure 6 plots this effective equation of state vs redshift for the two classes of models considered here. These plots show how $\omega_{\chi\,{\rm eff}}$ can easily take values less than -1 (the ‘phantom’ regime) despite $\omega_{\chi}\geq-1$ always being true. In particular $\omega_{\chi\text{eff}}<-1$ can hold right up to relatively small redshifts (around $z\sim 0.5$) in the case with a pure exponential dilaton potential for all coupling strengths, remaining in the phantom regime well into matter domination. When the dilaton has a local potential minimum into which to be trapped, there are much smaller crossings into apparently phantom territory as the dilaton oscillates during dark energy domination, but much deeper excursions at larger redshifts. The redshift at which this larger transition into the phantom regime occurs increases with dilaton-axion coupling strength.
Figure 6: Effective dilaton equations of state as a function of redshift for the case of an exponential dilaton potential described by (2.44) and (2.45) with $u_{1}=u_{2}=0$ on the left and trapped in a local quadratic minimum described by (2.44) and (2.45) on the right.
@@ Acknowledgements
We thank Elsa Teixeira for the useful numerical resources. This work evolved out of discussions at the Astroparticle Symposium at the Institut Pascal. CB, CvdB and ACD thank the Institut Pascal for their hospitality during the programme. MM is also grateful for the hospitality of Perimeter Institute where part of this work was carried out. AS is supported by the W.D. Collins Scholarship. CvdB is supported by the Lancaster–Sheffield Consortium for Fundamental Physics under STFC grant: ST/X000621/1. ACD is partially supported by the Science and Technology Facilities Council (STFC) through the STFC consolidated grant ST/T000694/1. CB’s research was partially supported by funds from the Natural Sciences and Engineering Research Council (NSERC) of Canada. MM is supported Kavli IPMU which was established by the World Premier International Research Center Initiative (WPI), MEXT, Japan. This work was also supported by a grant from the Simons Foundation (1034867, Dittrich). Research at the Perimeter Institute is supported in part by the Government of Canada through NSERC and by the Province of Ontario through MRI.